# Physical properties of elementary particles: Inertia and Interaction

### 09 Mar 2026 v1

@@ Abstract
Matter has two physical properties: Inertia and interaction. If we define the center of mass of an elementary particle in relation to its inertia, and a center of interaction in relation to its interactive properties, there are only two possibilities to describe this elementary particle: that both points are the same or that they are different. If they are the same, what we describe is the point particle model, while if we consider them to be different, what we obtain is the description of an elementary spinning particle. If the center of interaction or center of charge is moving at the speed of light, completely determines also the dynamics of the center of mass, and when quantizing this model satisfies Dirac’s equation. We obtain the classical description of the spinning Dirac particle.
@@ 1 Introduction
Since they are very small objects, we can accept that with respect to inertia there will exist a characteristic point called its center of mass (CM), and that with respect to its interaction properties there will exist at least one point that we will call the center of interaction or center of charge (CC).
@@
If we consider the boundary variables manifold given by $x\equiv(t,{\bi r}_{1},{\bi r}_{2})$ in terms of the position of two points, ${\bi r}_{1}$ and ${\bi r}_{2}$, two possible centers of two different interactions, then this manifold is not a homogeneous space of the Poincaré group. Given two sets of the above variables, a time translation transforms one time into the other. A space translation can transform one of the points, say ${\bi r}_{1}$, in its image, but it is impossible that the same space translation links the arbitrary initial and final positions of the other point ${\bi r}_{2}$. The classical elementary particle has to have a unique interaction center for all possible interactions. For quarks, the same point $\bi r$} represents the location of the center of electric charge, color charge and weak charge. Concerning gravitation, if we identify inertial mass with gravitational mass, it seems to suggest that the CM is the location of the gravitational interaction center, but this has to be consistent with a general theory of gravitation of spinning particles.
@@ 2 Constants of the motion of the free Dirac particle
The free Lagrangian of the Dirac particle is a function $L_{0}({\bi u},{\bi a},\boldsymbol{\omega})$ of the velocity and acceleration of the point $\bi r$}, moving at the speed of light $u=c$, and of the angular velocity of the comoving frame. This point satisfies a system of fourth order differential equations. Translation invariance implies that is independent of $t$ and $\bi r$}. For this particle, the Noether analysis of the constants of the motion [1], the invariance of $L_{0}$ under rotations defines the total angular momentum with respect to the origin of the observer frame as:
@@
is the spin with respect to the center of charge $\bi r$}. The different terms come from the Lagrangian as  [^1] In this work, three-dimensional expressions like ${\bi a}={\partial L_{0}}/{\partial{\bi b}}$, have to be interpreted as $a_{i}={\partial L_{0}}/{\partial b_{i}}$, $i=1,2,3$.:
${\bi p}=\frac{\partial L_{0}}{\partial{\bi u}}-\frac{d}{dt}\left(\frac{\partial L_{0}}{\partial{\bi a}}\right),\quad{\bi U}=\frac{\partial L_{0}}{\partial{\bi a}},\quad{\bi W}=\frac{\partial L_{0}}{\partial{\boldsymbol{\omega}}}.$
The linear momentum $\bi p$} is the conjugate momentum of the variables $\bi r$}, $\bi U$} is the conjugate momentum of the variables $\bi u$}, and finally $\bi W$} is the conjugate momentum of the orientation variables $\boldsymbol{\alpha}$}. The Hamiltonian is defined as usual as:
$H={\bi p}\cdot{\bi u}+{\bi U}\cdot{\bi a}+{\bi W}\cdot\boldsymbol{\omega}-L_{0}.$ (2)
Invariance of $L_{0}$ under pure Lorentz transformations defines the kinematical momentum with respect to the origin of the observer frame:
@@
Since $H$ and $\bi p$} are constants of the motion for the free particle, the time derivative of $\bi K$} leads to ${\bi p}=H{\bi v}/c^{2}$, where ${\bi v}=d{\bi q}/dt$ is the constant velocity of the point $\bi q$}, that can be interpreted as the center of mass (CM) of the Dirac particle. Because the invariant and constant of the motion $H^{2}-{\bi p}^{2}c^{2}=m^{2}c^{4}$ defines the mass of the particle, we arrive to
@@
$({\bi S}\times{\bi u})\times{\bi u}=H{\bi k}\times{\bi u},\quad-{\bi S}c^{2}=H({\bi r}-{\bi q})\times{\bi u},$ (7)
@@
The CM spin is conserved but the spin with respect to the CC satisfies the same dynamical equation than Dirac’s spin operator in the quantum case. It is the classical spin equivalent to Dirac’s spin operator. In the center of mass frame (${\bi q}={\bi v}=0$) and both spins take the same value. In the center of mass frame and for a free electron, both spins are constant point $\bi r$} satisfies the differential equation
@@
Figure 1: Motion of the CC of the electron $\bi r$} in the center of mass frame given in (11). The CM is located at the origin in this frame. The angular velocity $\boldsymbol{\omega}$, because ${\bi v}=0$ has no component along the velocity $\bi u$}, the trajectory is flat and has no torsion. The spin in the center of mass frame has the opposite direction to the angular velocity $\boldsymbol{\omega}$.
@@
$H{\bi u}-c^{2}{\bi p}-\frac{d{\bi S}}{dt}\times{\bi u}-{\bi S}\times\frac{d{\bi u}}{dt}=0.$
@@
The separation vector between the CM and CC, ${\bi q}-{\bi r}$, has the direction of the CC acceleration $\bi a$}. The CC performs a central motion around the CM.
@@ 2.1 Intrinsic properties of the Dirac particle
The observables $H$, $\bi p$}, the two spins $\bi S$}, ${\bi S}_{CM}$, the synchronous location of the CM $\bi q$} can be expressed in terms of the point $\bi r$} and their time derivatives, simultaneously.
@@ 3 The interaction Lagrangian
because the velocity is expressed as ${\bi u}=d{\bi r}/dt=\dot{\bi r}/\dot{t}$. It is also $\tau$-independent. Therefore $\widetilde{L}_{I}=L_{I}\dot{t}$, is a function of $\widetilde{L}_{I}(t,{\bi r},\dot{t},\dot{\bi r},)$, $\tau$-independent homogeneous function of first degree in terms of $\dot{t}$ and $\dot{\bi r}$. Euler’s theorem implies that
@@
and the vector function $\bi A$} is the partial derivative of the temporal Lagrangian $L_{I}$ with respect to $\bi u$}. For $A_{0}$ we make:
@@
As an example, the linear functions of $\bi u$}, ${\bi A}=f(t,{\bi r}){\bi u}$ and ${\bi A}={\bf g}(t,{\bi r})\times{\bi u}$, satisfy (19), but since $u^{2}=c^{2}$, the interaction Lagrangian reduces to the scalar function $L_{I}=V(t,{\bi r})$, independent of $\bi u$}. But in principle, we can find interaction Lagrangians more general than the minimal coupling.
@@ 4 Dynamical equations
The dynamical equations of the center of mass $\bi q$} and the center of charge $\bi r$} of the Dirac particle under any electromagnetic field obtained from the general Lagrangian $L=L_{0}+L_{em}$, where $L_{0}({\bi u},{\bi a},\boldsymbol{\omega})$ is the free Lagrangian and $L_{em}=-eA_{0}(t,{\bi r})+e{\bi u}\cdot{\bi A}(t,{\bi r})$, in terms of the scalar and vector potentials $A_{0}$ and $\bi A$} defined at time $t$ at the center of charge position $\bi r$}, respectively, are:
@@
where the external force ${\bi F}=e{\bi E}(t,{\bi r})+e{\bi u}\times{\bi B}(t,{\bi r})$, is the Lorentz force defined at the CC, and $m$ and $e$ are the mass and the electric charge of the electron, respectively. The two velocities satisfy the constraints $u=c$, $v<c$. Equation (20) is just $d{\bi p}/dt={\bi F}$, after taking the time derivative of ${\bi p}=\gamma(v)m{\bi v}$, and leaving on the left-hand side the CM acceleration $d{\bi v}/dt$. The equation (21) is equation (12) when writing on the left-hand side the CC acceleration $\bi a$}. The original fourth-order differential equations for the point $\bi r$} have been rewritten as a second-order system of differential equations for both points $\bi q$} and $\bi r$}, without specifying the Poincaré invariant free Lagrangian $L_{0}$, and with the minimal coupling interaction Lagrangian $L_{em}$.
@@
In the publication [5] we analyze the interaction of the Dirac particle with external uniform and oscillating electric and magnetic fields, we make a link to some Mathematica notebooks where the differential equations are solved numerically, and also to Wolfram’s Community groups [6] where some notebooks are included. In the work [7] we describe the interaction of the classical Dirac particle with a monochromatic circularly polarized plane wave, and in [8] we describe the interaction of two Dirac particles, where the Poincaré invariant interaction Lagrangian is velocity dependent, it is not of a minimal coupling structure but is a kind of instantaneous action at a distance of the two particles.

### 16 Mar 2026 v2

@@ Abstract
Matter has two physical properties: Inertia and interaction. If we define the center of mass of an elementary particle in relation to its inertia, and a center of interaction in relation to its interactive properties, there are only two possibilities to describe this elementary particle: that both points are the same or that they are different. If they are the same, what we describe is the point particle model, while if we consider them to be different, what we obtain is the description of an elementary spinning particle. If the center of interaction or center of charge is moving at the speed of light, completely determines also the dynamics of the center of mass, and when quantizing this model satisfies Dirac’s equation. We obtain the classical description of the spinning Dirac particle. The general analysis of the interaction Lagrangian, suggests a modification of the minimal coupling Lagrangian, for a possible classical description of the strong and weak interaction.
@@ 1 Introduction
Since they are very small objects, we can accept that with respect to inertia there will exist a characteristic point called its center of mass (CM), and that with respect to its interaction properties there will exist at least one point that we will call the center of interaction or center of charge (CC). With this assumption we consider that elementary particles are massive particles.
@@
If we consider the boundary variables manifold given by $x\equiv(t,{\bi r}_{1},{\bi r}_{2})$ in terms of the position of two points, ${\bi r}_{1}$ and ${\bi r}_{2}$, two possible centers of two different interactions, then this manifold is not a homogeneous space of the Poincaré group. Given two sets of the above variables, a time translation transforms one time into the other. A space translation can transform one of the points, say ${\bi r}_{1}$, in its image, but it is impossible that the same space translation links the arbitrary initial and final positions of the other point ${\bi r}_{2}$. The classical elementary particle has to have a unique interaction center for all possible interactions. For quarks, the same point $\bi r$} represents the location of the center of electric charge, color charge and weak charge. For the classical description of leptons, the point $\bi r$} represents the location of the electric and weak charge. Concerning gravitation, if we identify inertial mass with gravitational mass, it seems to suggest that the CM is the location of the gravitational interaction center, but this has to be consistent with a general theory of gravitation of spinning particles.
@@ 2 Constants of the motion of the free Dirac particle
The free Lagrangian of the Dirac particle is a function $L_{0}({\bi u},{\bi a},\boldsymbol{\omega})$ of the velocity and acceleration of the point $\bi r$}, moving at the speed of light $u=c$, and of the angular velocity of the comoving frame. It is translation invariant and, therefore, independent of $t$ and $\bi r$}. This point satisfies a system of fourth order differential equations. The generalized coordinates are $\bi r$}, $\bi u$} and $\boldsymbol{\alpha}$, the boundary variables with the time excluded. The dynamical equations of the $\bi r$} variables are:  [^1] In this work, three-dimensional expressions like ${\bi a}={\partial L_{0}}/{\partial{\bi b}}$, have to be interpreted as $a_{i}={\partial L_{0}}/{\partial b_{i}}$, $i=1,2,3$.
$\frac{\partial L_{0}}{\partial{\bi r}}-\frac{d}{dt}\left(\frac{\partial L_{0}}{\partial{\bi u}}\right)+\frac{d^{2}}{dt^{2}}\left(\frac{\partial L_{0}}{\partial{\bi a}}\right)=0,$
and because $L_{0}$ is space translation invariant, independent of $\bi r$}, the first term cancels out and the above fourth-order differential equation is
$\frac{d}{dt}\left[\frac{\partial L_{0}}{\partial{\bi u}}-\frac{d}{dt}\left(\frac{\partial L_{0}}{\partial{\bi a}}\right)\right]=-\frac{d{\bi p}}{dt}=0,$
and the conserved linear momentum, the canonical conjugate momentum of the $\bi r$} variables, is the term inside the squared brackets:
${\bi p}=\frac{\partial L_{0}}{\partial{\bi u}}-\frac{d}{dt}\left(\frac{\partial L_{0}}{\partial{\bi a}}\right).$ (1)
The dynamical equations of the orientation variables $\boldsymbol{\alpha}$, are:
$\frac{\partial L_{0}}{\partial{\boldsymbol{\alpha}}}-\frac{d}{dt}\left(\frac{\partial L_{0}}{\partial{d\boldsymbol{\alpha}/dt}}\right)=0,$
but the dependence of $L_{0}$ of $\boldsymbol{\alpha}$ and $d\boldsymbol{\alpha}/dt$ is through its dependence of the angular velocity $\boldsymbol{\omega}$, the above equation reduces to
$\frac{d{\bi W}}{dt}=\omega\times{\bi W},\quad{\rm where}\quad{\bi W}=\frac{\partial L_{0}}{\partial{\boldsymbol{\omega}}},$
and $\bi W$} is the canonical conjugate momentum of the orientation variables $\boldsymbol{\alpha}$}. The canonical conjugate momentum of the variables $\bi u$}, is finally,
${\bi U}=\frac{\partial L_{0}}{\partial{\bi a}}.$
The Hamiltonian is defined as usual as:
$H={\bi p}\cdot{\bi u}+{\bi U}\cdot{\bi a}+{\bi W}\cdot\boldsymbol{\omega}-L_{0},$ (2)
and is the conserved observable related to the time invariance of $L_{0}$.
The invariance of $L_{0}$ under rotations defines the total angular momentum with respect to the origin of the observer frame as [1]:
@@
is the spin with respect to the center of charge $\bi r$}. Please remark that the spin is related to the dependence of the Lagrangian of the acceleration $\bi a$}, ${\bi U}={\partial L_{0}}/{\partial{\bi a}}$, and the angular velocity $\boldsymbol{\omega}$, ${\bi W}={\partial L_{0}}/{\partial{\boldsymbol{\omega}}}$.
Invariance of $L_{0}$ under pure Lorentz transformations defines the kinematical momentum with respect to the origin of the observer frame [1]:
@@
Since $H$ and $\bi p$} are constants of the motion for the free particle, the time derivative of $\bi K$} leads to ${\bi p}=H{\bi v}/c^{2}$, where ${\bi v}=d{\bi q}/dt$ is the constant velocity of the point $\bi q$}. This point can be interpreted as the center of mass (CM) of the Dirac particle. Because the invariant and constant of the motion $H^{2}-{\bi p}^{2}c^{2}=m^{2}c^{4}$ defines the mass of the particle, we arrive to
@@
$({\bi S}\times{\bi u})\times{\bi u}=H{\bi k}\times{\bi u},\quad\rightarrow\quad-{\bi S}c^{2}=H({\bi r}-{\bi q})\times{\bi u},$ (8)
@@
It is expressed in terms of the separation of the two points $\bi r$} and $\bi q$}, and of the velocities of both points.
@@
It also depends on the separation of the two points $\bi r$} and $\bi q$}, and of the velocities of both points.
@@
The CM spin is conserved but the spin with respect to the CC satisfies the same dynamical equation than Dirac’s spin operator in the quantum case. It is the classical spin equivalent to Dirac’s spin operator. In the center of mass frame (${\bi q}={\bi v}=0$) and both spins take the same value. In the center of mass frame and for a free electron, both spins are constant, and point $\bi r$} satisfies the differential equation
@@
Figure 1: Motion of the CC of the electron $\bi r$} in the center of mass frame given in (12). The CM is located at the origin in this frame. The angular velocity $\boldsymbol{\omega}$, because ${\bi v}=0$ has no component along the velocity $\bi u$}, the trajectory is flat and has no torsion. If the CM velocity ${\bi v}\neq 0$ the angular velocity has a component along $\bi u$}, that produces the torsion of the trajectory of the $C$. The spin in the center of mass frame has the opposite direction to the angular velocity $\boldsymbol{\omega}$. Remark that the orientation of the spin with respect to the plane of the trajectory of point $\bi r$}, is opposite to the orientation of the angular momentum if the trajectory were the motion of a point mass.
@@
$H{\bi u}-c^{2}{\bi p}-\frac{d{\bi S}}{dt}\times{\bi u}-{\bi S}\times\frac{d{\bi u}}{dt}=0.$ (13)
@@
The separation vector between the CM and CC, ${\bi q}-{\bi r}$, has the direction of the CC acceleration $\bi a$} and is orthogonal to the CC velocity $\bi u$}. The CC performs a central motion at the constant velocity $c$, around the CM. All variables in the expression (14) are considered at the same time $t$, the time coordinate of the corresponding inertial observer.
@@ 2.1 Intrinsic properties of the Dirac particle
The observables $H$, $\bi p$}, the two spins $\bi S$}, ${\bi S}_{CM}$, the synchronous location of the CM $\bi q$}, can be expressed in terms of the point $\bi r$} and their time derivatives, simultaneously.
2.2 Dirac’s Hamiltonian
If in the above equation (13) we make the scalar product with $\bi u$} we get:
$H={\bi p}\cdot{\bi u}+\frac{1}{c^{2}}{\bi S}\cdot\left({\bi a}\times{\bi u}\right),$ (16)
This linear expression in terms of $H$ and $\bi p$}, represents the classical equivalent of Dirac’s Hamiltonian. The energy of the Dirac particle has two terms: the first ${\bi p}\cdot{\bi u}$, represents the translation energy, depends on the mass and is related to the inertia of the particle. The second, ${\bi S}\cdot\left({\bi a}\times{\bi u}\right)/c^{2}$, is positive definite and non-vanishing, represents the rotation energy and depends on the other intrinsic property of the Dirac particle, the spin. Because $\bi p$} is the canonical conjugate momentum of the variable $\bi r$}, we obtain the Hamilton equation $\partial H/\partial{\bi p}={\bi u}$.
@@ 3 The interaction Lagrangian
because the velocity is expressed as ${\bi u}=d{\bi r}/dt=\dot{\bi r}/\dot{t}$. It is also $\tau$-independent. Therefore $\widetilde{L}_{I}=L_{I}\dot{t}$, is a function of $\widetilde{L}_{I}(t,{\bi r},\dot{t},\dot{\bi r})$, $\tau$-independent homogeneous function of first degree in terms of $\dot{t}$ and $\dot{\bi r}$. Euler’s theorem implies that
@@
and the vector function ${\bi A}=\partial{L}_{I}/\partial{\bi u}$, is the partial derivative of the temporal Lagrangian $L_{I}$ with respect to $\bi u$}. For $A_{0}$ we make:
@@
As an example, the linear functions of $\bi u$}, ${\bi A}=f(t,{\bi r}){\bi u}$ and ${\bi A}={\bf g}(t,{\bi r})\times{\bi u}$, satisfy (22), but since $u^{2}=c^{2}$, the interaction Lagrangian reduces to the scalar function $L_{I}=V(t,{\bi r})$, independent of $\bi u$}. But in principle, we can look for interaction Lagrangians more general than the minimal coupling to describe other classical interactions.
Since $A_{0}$ is independent of $\bi u$}, from (21) we obtain $({\partial A_{j}}/{\partial u_{k}})u_{k}=0$, so that $\bi A$} is a homogeneous function of $\bi u$} of zero degree. It is a function of ${\bi u}/u$, and thus this possible interaction Lagrangian must depend, in addition to the time $t$ and position $\bi r$}, on the two compact variables, the zenithal angle $\theta$ and the azimuthal angle $\phi$, of the non-vanishing CC velocity $\bi u$}. This more general interaction Lagrangian must be dependent of the orientation of the CC velocity of the particle.
@@ 4 Dynamical equations
The dynamical equations of the center of mass $\bi q$} and the center of charge $\bi r$} of the Dirac particle under any electromagnetic field are obtained from the general Lagrangian $L=L_{0}+L_{em}$, where $L_{0}({\bi u},{\bi a},\boldsymbol{\omega})$ is the free Lagrangian. The interaction Lagrangian $L_{em}=-eA_{0}(t,{\bi r})+e{\bi u}\cdot{\bi A}(t,{\bi r})$, is written in terms of the electric charge $e$ and the scalar and vector potentials $A_{0}$ and $\bi A$} defined at time $t$ at the center of charge position $\bi r$}. Euler-Lagrange equations give rise to the first order system:
@@
where the external force ${\bi F}=e{\bi E}(t,{\bi r})+e{\bi u}\times{\bi B}(t,{\bi r})$, is the Lorentz force defined at the CC, and $m$ and $e$ are the mass and the electric charge of the electron, respectively. The two velocities satisfy the constraints $u=c$, $v<c$. Equation (23) is just the Euler-Lagrange equation of the $\bi r$} variables $d{\bi p}/dt={\bi F}$, after taking the time derivative of ${\bi p}=\gamma(v)m{\bi v}$, and leaving on the left-hand side the CM acceleration $d{\bi v}/dt$. The equation (24) is equation (14) when writing on the left-hand side the CC acceleration $\bi a$}. The original fourth-order differential equations for the point $\bi r$} have been rewritten as a second-order system of differential equations for both points $\bi q$} and $\bi r$}, or as the first order system (23) and (24) for the variables $\bi q$}, $\bi v$}, $\bi r$} and $\bi u$}, more suitable for a numerical integration, without specifying the Poincaré invariant free Lagrangian $L_{0}$, and with the minimal coupling interaction Lagrangian $L_{em}$.
@@
In the publication [5] we analyze the interaction of the Dirac particle with external uniform and oscillating electric and magnetic fields, we make a link to some Mathematica notebooks where the differential equations are solved numerically, and also to Wolfram’s Community groups [6] where some notebooks are included. In the work [7] we describe the interaction of the classical Dirac particle with a monochromatic circularly polarized plane wave, and in [8] we describe the interaction of two Dirac particles, where the Poincaré invariant interaction Lagrangian is velocity dependent, it is not of a minimal coupling structure but is a kind of instantaneous action at a distance of the two particles, and depends on the relative orientations of the CC velocities of both particles, as suggested in Section 3 when describing the general interaction Lagrangian. Because electrons are leptons, a possible classical weak interaction description of leptons, should have a weak interaction Lagrangian that depends on the orientation of the CC velocity.
5 Radiation Reaction
We have considered that elementary particles have inertia and therefore they are massive particles. The main characteristic is that they have a CC and CM as two different points. In [9] we have shown that in the case of the electromagnetic interaction the radiation occurs because the energy expended by the field is the work of the external Lorentz force along the CC trajectory, while the change of the mechanical energy of the particle is the work along the CM trajectory. If these two works are different, energy conservation of the closed system particle+field, implies that there is a continuous transfer of electromagnetic energy, linear momentum and angular momentum between the particle and the field. Preliminary calculations of the transfer of energy and linear momentum, show that the amount of energy equals the value of the linear momentum transfer and this can be interpreted as if a massless electromagnetic quantum of energy and linear momentum has been transferred to the field. To consider that this transfer corresponds to a photon we have to compute the transfer of angular momentum. Because the transfer is continuous, this will correspond to a photon if after some time $\Delta t$, the total continuous transfer of angular momentum amounts $\hbar$, the angular momentum of a photon.
The requirement that the value of the spin for the center of mass observer has led to modify the dynamical equation (23) $d{\bi p}/dt={\bi F}$, by introducing a breaking or accelerating term along the CM velocity $\bi v$},
$\frac{d{\bi p}}{dt}={\bi F}-\gamma^{2}\left[({\bi u}-{\bi v})\cdot{\bi F}\right]{\bi v},$
term that is related to the difference of the work per unit time, of the external Lorentz force $\bi F$} along the CC and CM trajectories.
An accelerated electron by an external Lorentz force, needs some time $\Delta t$, to interchange energy linear momentum and also a unit $\hbar$ of angular momentum with the external field. If this happens to be correct, the interaction between the field and the particle corresponds to the creation of a quantum of electromagnetic energy, linear momentum and of angular momentum 1, in natural units. With the above consideration that elementary particles are massive particles, the created photon is not an elementary particle. We cannot define its center of mass, but corresponds to an uncharged massless object that moves at the speed of light along a straight line. The electromagnetic interaction implies the assumed interchange of massless bosons of spin 1, as derived from Quantum Electrodynamics, that can be created and annihilated during the process of interaction. The photon will be the result of the electromagnetic interaction.
What follows is a conjecture. Let us assume that with the freedom we have to determine another possible interaction Lagrangian, we were able to obtain a classical description of other interactions. The CC will be the same interacting point as the center of electric charge of the Dirac particle, and being different than the CM, there will be a difference between the external work and the increase of the mechanical energy of the particle for the new interaction. If it happens that this interchange of mechanical properties between the particle and the field produces new particles, they will be bosons because the angular momentum of the Dirac particle has to remain the same. If they are massless, they must be moving at the speed of light and must be bosons because the spin of leptons and quarks is $1/2$.
We have to obtain first a new non-minimal coupling interaction Lagrangian, that predicts a short-range interacting force, but the results of quantum chromodynamics and weak interactions predict, in the quantum analysis, the existence of intermediate bosons, massless gluons in the strong interaction and very massive bosons for the weak interaction. The existence of massive bosons has also been considered in the general formalism [1]. Spinning particles with a CC different than the CM, such that the CC is moving faster than light can also be described in this formalism. The CM is moving at a velocity $v<c$, but when quantizing these models, they only produce integer spin particles. Faster than light motions of the CC of an elementary particle only quantize with integer values. If this happens to be correct, the $Z^{0}$ and $W^{\pm}$ massive bosons of the weak interaction, will correspond, from the classical point of view, to massive particles with a CM that can be at rest, but whose CC is moving faster than light.